Let B be a 2-dimensional ring with an action of a group G. The combinatorial data associated with the desingularization of the ring of invariants BG produces a symmetric semi-definite negative integer quadratic form, and properties of this form will be described in the wild case, that is, when the residual characteristic of B is not coprime to the order of G. This is intended to be a friendly introduction to this subject, and motivations and examples will be included.